OpenStudy (anonymous):

Megan and Sam together take 12 minutes to wash a car. Washing the car by herself Megan took 21 minutes. How long would it take Sam to wash the car?

OpenStudy (lgbasallote):

let x = sam's time y = megan's ime use the formula \[\frac tx + \frac ty = 1\] where t is total time now substitute \[\frac{12}{x} + \frac{12}{21} = 1\] now we solve for x.

OpenStudy (lgbasallote):

do you need more help?

OpenStudy (anonymous):

Can you solve for x and show me the steps?

OpenStudy (lgbasallote):

subtract both sides by \(\frac{12}{21}\) \[\frac{12}{x} = 1 - \frac{12}{21}\] \[\frac{12}{x} = \frac{21}{21} - \frac{12}{21}\] \[\frac{12}{x} = \frac{9}{21}\] multiply both sides by x \[12 = \frac{9}{21} x\] multiply both sides by 21 \[12 \times 21 = 9x\] divide both sides by 9 \[\frac{12 \times 21}{9} = x\] simplify

OpenStudy (lgbasallote):

i need to go now so i hope you got that!